# The Ginsburg–Sands theorem and computability theory

### v1

@@ 4. General spaces
That (1) implies (2) follows readily from the proof of the Ginsburg–Sands theorem given in Section 2. More precisely, given a CSC space with its closure relation, we can construct the equivalence relation $\sim$ and the partial ordering $\leq_{X}$ defined in the proof, and argue as before to obtain an infinite subspace which is either indiscrete, has the initial segment or final segment topologies, or is $T_{1}$. For the reverse implication, assume $\mathsf{wGS}^{\operatorname{cl}}$ and let $(P,\leq_{P})$ be any infinite partial order. Assume this partial ordering has no infinite antichain; we will construct an infinite chain. (In fact, our chain will either be an infinite ascending sequence under $\leq_{P}$, or an infinite descending sequence.) For each $p\in P$, define $V_{p}=\{q\in P:p\leq_{P}q\}$ and let $\langle P,\mathcal{U},k\rangle$ be the CSC space generated by $\langle V_{p}:p\in P\rangle$. We claim that, for $p,q\in P$, $p$ is in the closure of $q$ in this space if and only if $p\leq_{P}q$. Indeed, if $p$ is in the closure of $q$ then in particular $q$ must belong to $V_{p}$. So $p\leq_{P}q$. Suppose $p$ is not in the closure of $q$. Then we can fix a $U\in\mathcal{U}$ containing $p$ but not $q$. By Proposition 3.3 (2), there is an $r\in P$ such that $V_{r}$ contains $p$ but not $q$, meaning $r\leq_{P}p$ but $r\nleq_{P}q$. So $p\nleq_{P}q$. This proves the claim, and implies that the closure relation exists. Call it $\operatorname{cl}$. Now apply $\mathsf{wGS}^{\operatorname{cl}}$ to obtain an infinite subspace $Y$ of $\langle P,\mathcal{U},k\rangle$. By assumption, $Y$ is not an infinite antichain under $\leq_{P}$. Thus we can fix some $p<_{P}q$ in $Y$. It follows that $p\notin U_{q}$, so $Y$ cannot be indiscrete as a subspace. It also cannot be $T_{1}$, because if it were we would have $p\notin\operatorname{cl}(q)$ and hence $p\nleq_{P}q$. Thus, $Y$ either has the initial segment or final segment topology. Consider first the case that it is the former. Then there is a bijection $h:\mathbb{N}\to Y$ such that the basic open sets of $Y$ in the subspace topology are precisely $\{h(i):i\leq j\}$ for some $j\in\mathbb{N}$. We claim that $h(j)<_{P}h(i)$ for all $i<j$, which proves that $Y$ is a descending sequence under $\leq_{P}$. Suppose not. Then there exist $p,q\in Y$ such that $h^{-1}(q)<h^{-1}(p)$ and $p\nleq_{P}q$. Then $p\notin\operatorname{cl}(q)$, so there exists $U\in\mathcal{U}$ containing $p$ but not $q$. Fix $j$ so that $U\cap Y=\{h(i):i\leq j\}$. Then necessarily $h^{-1}(p)\leq j$, but this means that also $h^{-1}(q)\leq j$ and hence that $q\in U$, a contradiction. The case where $Y$ has the final segment topology is symmetric, and yields an ascending sequence under $\leq_{P}$. ∎
@@
We argue in $\mathsf{RCA}_{0}$. To prove that (2) implies (1), assume $\mathsf{GST}_{1}$ and let a CSC space $\langle X,\mathcal{U},k\rangle$ with closure relation $\operatorname{cl}$ be given. By Corollary 6.10 below, $\mathsf{GST}_{1}$ implies $\mathsf{RT}^{2}_{2}$ and therefore $\mathsf{CAC}$. Hence, by Theorem 4.7, $\mathsf{GST}_{1}$ implies $\mathsf{wGS}^{\operatorname{cl}}$. So let $Y\subseteq X$ be an infinite subspace obtained by applying $\mathsf{wGS}^{\operatorname{cl}}$ to $X$. If $Y$ is indiscrete, or homeomorphic to the initial segment or final segment topologies, then $Y$ is a solution to $X$ as an instance of $\mathsf{GS}^{\operatorname{cl}}$. If $Y$ is $T_{1}$, then we can apply $\mathsf{GST}_{1}$ to $Y$ to obtain a subspace $Z$ of $Y$ (and so, of $X$) that is discrete or has the cofinite topology. To prove that (1) implies (2), assume $\mathsf{GS}^{\operatorname{cl}}$ and let $\langle X,\mathcal{U},k\rangle$ be a $T_{1}$ CSC space. The closure relation on $X$ is then trivial, since $y\in\operatorname{cl}(x)$ if and only if $x=y$. We can thus apply $\mathsf{GS}^{\operatorname{cl}}$ to $X$ to obtain an infinite subspace $Y\subseteq X$, and since $X$ is $T_{1}$ this subspace cannot be indiscrete, nor be homeomorphic to either the initial segment or final segment topologies. It follows that $Y$ is discrete or has the cofinite topology, as needed. ∎
@@ 5. Hausdorff spaces
Since $c$ is $\Delta^{0}_{1}$ definable, to prove (1) it suffices to prove that it is total. Fix $x\neq y$ in $X$ and $s\in\mathbb{N}$. Since $X$ is Hausdorff, there exist $U_{m},U_{n}$ such that $x\in U_{m}$, $y\in U_{n}$, and $U_{m}\cap U_{n}=\emptyset$. In particular, the set of all pairs $\langle m,n\rangle$ such that $x\in U_{m}\smallsetminus U_{n}$, $y\in U_{n}\smallsetminus U_{m}$, and $(U_{m}\operatorname{\upharpoonright}s)\cap(U_{n}\operatorname{\upharpoonright}s)=\emptyset$ is nonempty. Now $\mathsf{I}\Sigma^{0}_{1}$ suffices to find the least element $\langle m,n\rangle$ of this set, and we have $c(x,y,s)=\langle m,n\rangle$. To prove (2), we note that the least $\langle m,n\rangle$ such that $x\in U_{m}$, $y\in U_{n}$, and $U_{m}\cap U_{n}=\emptyset$ exists by $\mathsf{I}\Pi^{0}_{1}$. Thus, for each $\langle i,j\rangle<\langle m,n\rangle$ there exists a $t$ such that if $x\in U_{i}\smallsetminus U_{j}$ and $y\in U_{j}\smallsetminus U_{i}$ then $t\in U_{i}\cap U_{j}$. By $\mathsf{B}\Sigma^{0}_{0}$ in $\mathsf{RCA}_{0}$ we can find an $s_{0}$ such that for each $\langle i,j\rangle<\langle m,n\rangle$ there exists such a $t$ below $s_{0}$. Then $c(x,y,s)=\langle m,n\rangle$ for all $s\geq s_{0}$. ∎
@@
We argue in $\mathsf{RCA}_{0}$. Let $(X,\mathcal{U},k)$ be an infinite Hausdorff CSC space with $\mathcal{U}=\langle U_{n}:n\in\mathbb{N}\rangle$. If $X$ has no limit point, then $X$ is already discrete. So, suppose $X$ has a limit point, $p$. We define a sequence $x_{0}<x_{1}<\cdots\in X$ by primitive recursion such that $x_{i}\neq p$ for all $i$. Let $x_{0}$ be any element of $X$ different from $p$. Now fix $s$, and suppose we have defined $x_{i}$ for all $i\leq s$. For each $i\leq s$, let $\langle m_{i,s},n_{i,s}\rangle=c(p,x_{i},s)$ and let $m_{s}=\overline{k}(p,\{m_{i,s}:i\leq s\})$, where $\overline{k}$ and $c$ are as in Lemmas 5.1 and 3.2. In particular, $p\in U_{m_{i,s}}$ for all $i\leq s$, so $p\in U_{m_{s}}\subseteq\bigcap_{i\leq s}U_{m_{i,s}}$. Since $p$ is a limit point, $U_{m_{s}}$ is infinite, and we can let $x_{s+1}$ be any element of this set larger than $x_{s}$ and different from $p$. Let $Y=\{x_{0},x_{1},\ldots\}$. We claim that $Y$ has the discrete topology. To see this, fix $i$, and let $\langle m,n\rangle$ be least such that $p\in U_{m}$, $x_{i}\in U_{n}$, and $U_{m}\cap U_{n}=\emptyset$. By Lemma 5.1 (2), there is an $s_{0}>i$ such that $c(p,x_{i},s)=\langle m,n\rangle$ for all $s\geq s_{0}$. For all $s\geq s_{0}$, the point $x_{s}$ is chosen in $U_{m,s}\subseteq U_{m_{i,s}}=U_{m}$, which means that $U_{n}\cap Y\subseteq\{x_{j}:j<s\}$. Thus, $U_{n}\cap Y$ is a finite basic open set in $Y$ containing $x_{i}$. Since $Y$ is Hausdorff by Proposition 3.18, and so in particular $T_{1}$, it follows by Proposition 3.7 that $\{x_{i}\}$ is open in $Y$. ∎
@@
Let $(X,\mathcal{U},k)$ be an infinite effectively Hausdorff CSC space which is not discrete. Then $X$ has a limit point, $p$. This means that for every $U\in\mathcal{U}$ containig $p$, there exist infinitely many $x\in U$ different from $p$. Let $\mathcal{U}=\langle U_{n}:n\in\mathbb{N}\rangle$, and let $e:X\times X\to\mathbb{N}$ be such that $e(x,y)=\langle m,n\rangle$ for all distinct $x,y\in X$, where $x\in U_{m}$, $y\in U_{n}$, and $U_{m}\cap U_{n}=\emptyset$. We construct sequences $x_{0}<x_{1}<\cdots\in X$, $n_{0},n_{1},\ldots\in\mathbb{N}$, and $m_{0},m_{1},\ldots\in\mathbb{N}$ such that $U_{m_{0}}\supseteq U_{m_{1}}\supseteq\cdots$, and for all $i$ and $j$: $x_{j}\in U_{n_{i}}$ if and only if $i=j$; $p\in U_{m_{i}}$; and $U_{m_{i}}\cap U_{n_{i}}=\emptyset$. Then $Y=\{x_{i}:i\in\mathbb{N}\}$ is an effectively discrete subspace of $X$ since $U_{n_{i}}\cap Y=\{x_{i}\}$ for all $i$. Let $x_{0}$ be any element of $X$ different from $p$, and let $n_{0}$ and $m_{0}$ be such that $e(p,x_{0})=\langle m_{0},n_{0}\rangle$. Now fix $s$, and suppose we have defined $x_{i}$, $n_{i}$, and $m_{i}$ for all $i\leq s$. Since $p\in\bigcap_{i\leq s}U_{m_{i}}$ we have that $p\in U_{\overline{k}(p,\{m_{i}:i\leq s\})}\subseteq\bigcap_{i\leq s}U_{m_{i}}$. And since $p$ is a limit point, we can let $x_{s+1}$ be the least element of $U_{\overline{k}(p,\{m_{i}:i\leq s\})}$ larger than $x_{s}$ and different from $p$. Say $e(p,x_{s})=\langle m,n\rangle$. Then $p\in\bigcap_{i\leq s}U_{m_{i}}\cap U_{m}$ and $x_{s+1}\in U_{m_{s}}\cap U_{n}$, so $p\in U_{\overline{k}(p,\{m_{i}:i\leq s\}\cup\{m\})}\subseteq\bigcap_{i\leq s}U_{m_{i}}\cap U_{m}$, and $x_{s+1}\in U_{\overline{k}(x_{i},\{m_{s},n\})}\subseteq U_{m_{s}}\cap U_{n}$. Let $m_{s+1}=\overline{k}(p,\{m_{i}:i\leq s\}\cup\{m\})$ and $n_{s+1}=\overline{k}(x_{s+1},\{m_{s},n\})$. So, $x_{s+1}\in U_{n_{s+1}}$, $p\in U_{m_{s+1}}$, $U_{m_{s+1}}\subseteq U_{m_{s}}$, and $U_{m_{s+1}}\cap U_{n_{s+1}}=\emptyset$. Fix any $i\leq s$. We have $x_{s+1}\in U_{n_{s+1}}\subseteq U_{m_{s}}$, while $x_{i}\in U_{n_{i}}$ which is disjoint from $U_{m_{i}}$ and hence from $U_{m_{s}}\subseteq U_{m_{i}}$. Thus, $x_{i}\notin U_{n_{s+1}}$ and $x_{s+1}\notin U_{n_{i}}$, as needed. ∎
@@
We construct a computable sequence $\langle V_{n}:n\in\omega\rangle$ of subsets of $\omega$ and then generate a computable CSC space $\langle\omega,\mathcal{U},k\rangle$ using Proposition 3.3. Initially, we enumerate $x$ into $V_{2\langle x,y\rangle}$ and $V_{2\langle x,y\rangle+1}$ for all $y$. We then enumerate more elements into the $V_{n}$ in stages, maintaining throughout that $V_{2\langle x,y\rangle}\cap V_{2\langle y,x\rangle}=\emptyset$ for all $x\neq y$. By Proposition 3.3 (2), we can compute the indices $m$ and $n$ of $V_{2\langle x,y\rangle}$ and $V_{2\langle y,x\rangle}$ in $\mathcal{U}$ uniformly computably from $x$ and $y$. Hence, the function $e:X\times X\to\mathbb{N}$ mapping $\langle x,y\rangle$ to $\langle m,n\rangle$ is computable. Since $x\in V_{2\langle x,y\rangle}$ and $y\in V_{2\langle y,x\rangle}$, this $e$ will witness that $\langle\omega,\mathcal{U},k\rangle$ is effectively Hausdorff. We aim to additionally satisfy the following requirements for all $x,e,u\in\omega$:
@@
$\mathcal{R}_{e,u}$ ::: if $\Phi_{e}$ is the characteristic function of an infinite set $Y$ and $\Phi_{u}$ is total on $Y$, then there exists $x\in Y$ and $n_{0},\ldots,n_{r}\in\omega$ such that $\Phi_{u}(x)\operatorname{\downarrow}=\langle n_{0},\ldots,n_{n}\rangle$ but $\bigcap_{i\leq r}V_{n_{i}}\neq\{x\}$.
We call these the $\mathcal{D}$-requirements and $\mathcal{R}$-requirements for short. The $\mathcal{D}$-requirements thus ensure that our CSC space is discrete, while the $\mathcal{R}$-requirements ensure that it has no infinite computable effectively discrete subspace. At each stage, at most one requirement will act. If a $\mathcal{D}$-requirement acts it claims an index $n\in\omega$, and if an $\mathcal{R}$-requirement acts it claims finitely many indices $n_{0},\ldots,n_{r}\in\omega$. A requirement may have its claims cancelled at a later stage, and then make new claims later still, as we describe below. We proceed to the construction. We assign a priority ordering to the requirements in order type $\omega$, as usual. For notational convenience, for $n=2\langle x,y\rangle$ we let $\overline{n}$ denote $2\langle y,x\rangle$. Suppose we are at stage $s\in\omega$. Say $\mathcal{D}_{x}$ requires attention at stage $s$ if it does not currently have a claim. Say $\mathcal{R}_{e,u}$ requires attention at stage $s$ if it does not currently have any claims, and if there exist $x<s$ and $\langle n_{0},\ldots,n_{r}\rangle<s$ as follows:
@@
  • no $n_{i}$ is claimed by any stronger requirement;

### v2

@@ 4. General spaces
That (1) implies (2) follows readily from the proof of the Ginsburg–Sands theorem given in Section 2. More precisely, given a CSC space with its closure relation, we can construct the equivalence relation $\sim$ and the partial ordering $\leq_{X}$ defined in the proof, and argue as before to obtain an infinite subspace which is either indiscrete, has the initial segment or final segment topologies, or is $T_{1}$.
For the reverse implication, assume $\mathsf{wGS}^{\operatorname{cl}}$ and let $(P,\leq_{P})$ be any infinite partial order. Assume this partial ordering has no infinite antichain; we will construct an infinite chain. (In fact, our chain will either be an infinite ascending sequence under $\leq_{P}$, or an infinite descending sequence.)
For each $p\in P$, define $V_{p}=\{q\in P:p\leq_{P}q\}$ and let $\langle P,\mathcal{U},k\rangle$ be the CSC space generated by $\langle V_{p}:p\in P\rangle$. We claim that, for $p,q\in P$, $p$ is in the closure of $q$ in this space if and only if $p\leq_{P}q$. Indeed, if $p$ is in the closure of $q$ then in particular $q$ must belong to $V_{p}$. So $p\leq_{P}q$. Suppose $p$ is not in the closure of $q$. Then we can fix a $U\in\mathcal{U}$ containing $p$ but not $q$. By Proposition 3.3 (2), there is an $r\in P$ such that $V_{r}$ contains $p$ but not $q$, meaning $r\leq_{P}p$ but $r\nleq_{P}q$. So $p\nleq_{P}q$. This proves the claim, and implies that the closure relation exists. Call it $\operatorname{cl}$.
Now apply $\mathsf{wGS}^{\operatorname{cl}}$ to obtain an infinite subspace $Y$ of $\langle P,\mathcal{U},k\rangle$. By assumption, $Y$ is not an infinite antichain under $\leq_{P}$. Thus we can fix some $p<_{P}q$ in $Y$. It follows that $p\notin U_{q}$, so $Y$ cannot be indiscrete as a subspace. It also cannot be $T_{1}$, because if it were we would have $p\notin\operatorname{cl}(q)$ and hence $p\nleq_{P}q$. Thus, $Y$ either has the initial segment or final segment topology. Consider first the case that it is the former. Then there is a bijection $h:\mathbb{N}\to Y$ such that the basic open sets of $Y$ in the subspace topology are precisely $\{h(i):i\leq j\}$ for some $j\in\mathbb{N}$. We claim that $h(j)<_{P}h(i)$ for all $i<j$, which proves that $Y$ is a descending sequence under $\leq_{P}$. Suppose not. Then there exist $p,q\in Y$ such that $h^{-1}(q)<h^{-1}(p)$ and $p\nleq_{P}q$. Then $p\notin\operatorname{cl}(q)$, so there exists $U\in\mathcal{U}$ containing $p$ but not $q$. Fix $j$ so that $U\cap Y=\{h(i):i\leq j\}$. Then necessarily $h^{-1}(p)\leq j$, but this means that also $h^{-1}(q)\leq j$ and hence that $q\in U$, a contradiction. The case where $Y$ has the final segment topology is symmetric, and yields an ascending sequence under $\leq_{P}$. ∎
@@
We argue in $\mathsf{RCA}_{0}$. To prove that (1) implies (2), assume $\mathsf{GST}_{1}$ and let a CSC space $\langle X,\mathcal{U},k\rangle$ with closure relation $\operatorname{cl}$ be given. By Corollary 6.10 below, $\mathsf{GST}_{1}$ implies $\mathsf{RT}^{2}_{2}$ and therefore $\mathsf{CAC}$. Hence, by Theorem 4.7, $\mathsf{GST}_{1}$ implies $\mathsf{wGS}^{\operatorname{cl}}$. So let $Y\subseteq X$ be an infinite subspace obtained by applying $\mathsf{wGS}^{\operatorname{cl}}$ to $X$. If $Y$ is indiscrete, or homeomorphic to the initial segment or final segment topologies, then $Y$ is a solution to $X$ as an instance of $\mathsf{GS}^{\operatorname{cl}}$. If $Y$ is $T_{1}$, then we can apply $\mathsf{GST}_{1}$ to $Y$ to obtain a subspace $Z$ of $Y$ (and so, of $X$) that is discrete or has the cofinite topology.
To prove that (2) implies (1), assume $\mathsf{GS}^{\operatorname{cl}}$ and let $\langle X,\mathcal{U},k\rangle$ be a $T_{1}$ CSC space. The closure relation on $X$ is then trivial, since $y\in\operatorname{cl}(x)$ if and only if $x=y$. We can thus apply $\mathsf{GS}^{\operatorname{cl}}$ to $X$ to obtain an infinite subspace $Y\subseteq X$, and since $X$ is $T_{1}$ this subspace cannot be indiscrete, nor be homeomorphic to either the initial segment or final segment topologies. It follows that $Y$ is discrete or has the cofinite topology, as needed. ∎
@@ 5. Hausdorff spaces
Since $c$ is $\Delta^{0}_{1}$ definable, to prove (1) it suffices to prove that it is total. Fix $x\neq y$ in $X$ and $s\in\mathbb{N}$. Since $X$ is Hausdorff, there exist $U_{m},U_{n}$ such that $x\in U_{m}$, $y\in U_{n}$, and $U_{m}\cap U_{n}=\emptyset$. In particular, the set of all pairs $\langle m,n\rangle$ such that $x\in U_{m}\smallsetminus U_{n}$, $y\in U_{n}\smallsetminus U_{m}$, and $(U_{m}\operatorname{\upharpoonright}s)\cap(U_{n}\operatorname{\upharpoonright}s)=\emptyset$ is nonempty. Now $\mathsf{I}\Sigma^{0}_{1}$ suffices to find the least element $\langle m,n\rangle$ of this set, and we have $c(x,y,s)=\langle m,n\rangle$.
To prove (2), we note that the least $\langle m,n\rangle$ such that $x\in U_{m}$, $y\in U_{n}$, and $U_{m}\cap U_{n}=\emptyset$ exists by $\mathsf{I}\Pi^{0}_{1}$. Thus, for each $\langle i,j\rangle<\langle m,n\rangle$ there exists a $t$ such that if $x\in U_{i}\smallsetminus U_{j}$ and $y\in U_{j}\smallsetminus U_{i}$ then $t\in U_{i}\cap U_{j}$. By $\mathsf{B}\Sigma^{0}_{0}$ in $\mathsf{RCA}_{0}$ we can find an $s_{0}$ such that for each $\langle i,j\rangle<\langle m,n\rangle$ there exists such a $t$ below $s_{0}$. Then $c(x,y,s)=\langle m,n\rangle$ for all $s\geq s_{0}$. ∎
@@
We argue in $\mathsf{RCA}_{0}$. Let $(X,\mathcal{U},k)$ be an infinite Hausdorff CSC space with $\mathcal{U}=\langle U_{n}:n\in\mathbb{N}\rangle$. If $X$ has no limit point, then $X$ is already discrete. So, suppose $X$ has a limit point, $p$. We define a sequence $x_{0}<x_{1}<\cdots\in X$ by primitive recursion such that $x_{i}\neq p$ for all $i$. Let $x_{0}$ be any element of $X$ different from $p$. Now fix $s$, and suppose we have defined $x_{i}$ for all $i\leq s$. For each $i\leq s$, let $\langle m_{i,s},n_{i,s}\rangle=c(p,x_{i},s)$ and let $m_{s}=\overline{k}(p,\{m_{i,s}:i\leq s\})$, where $\overline{k}$ and $c$ are as in Lemmas 5.1 and 3.2. In particular, $p\in U_{m_{i,s}}$ for all $i\leq s$, so $p\in U_{m_{s}}\subseteq\bigcap_{i\leq s}U_{m_{i,s}}$. Since $p$ is a limit point, $U_{m_{s}}$ is infinite, and we can let $x_{s+1}$ be any element of this set larger than $x_{s}$ and different from $p$.
Let $Y=\{x_{0},x_{1},\ldots\}$. We claim that $Y$ has the discrete topology. To see this, fix $i$, and let $\langle m,n\rangle$ be least such that $p\in U_{m}$, $x_{i}\in U_{n}$, and $U_{m}\cap U_{n}=\emptyset$. By Lemma 5.1 (2), there is an $s_{0}>i$ such that $c(p,x_{i},s)=\langle m,n\rangle$ for all $s\geq s_{0}$. For all $s\geq s_{0}$, the point $x_{s}$ is chosen in $U_{m,s}\subseteq U_{m_{i,s}}=U_{m}$, which means that $U_{n}\cap Y\subseteq\{x_{j}:j<s\}$. Thus, $U_{n}\cap Y$ is a finite basic open set in $Y$ containing $x_{i}$. Since $Y$ is Hausdorff by Proposition 3.18, and so in particular $T_{1}$, it follows by Proposition 3.7 that $\{x_{i}\}$ is open in $Y$. ∎
@@
Let $(X,\mathcal{U},k)$ be an infinite effectively Hausdorff CSC space which is not discrete. Then $X$ has a limit point, $p$. This means that for every $U\in\mathcal{U}$ containig $p$, there exist infinitely many $x\in U$ different from $p$. Let $\mathcal{U}=\langle U_{n}:n\in\mathbb{N}\rangle$, and let $e:X\times X\to\mathbb{N}$ be such that $e(x,y)=\langle m,n\rangle$ for all distinct $x,y\in X$, where $x\in U_{m}$, $y\in U_{n}$, and $U_{m}\cap U_{n}=\emptyset$. We construct sequences $x_{0}<x_{1}<\cdots\in X$, $n_{0},n_{1},\ldots\in\mathbb{N}$, and $m_{0},m_{1},\ldots\in\mathbb{N}$ such that $U_{m_{0}}\supseteq U_{m_{1}}\supseteq\cdots$, and for all $i$ and $j$: $x_{j}\in U_{n_{i}}$ if and only if $i=j$; $p\in U_{m_{i}}$; and $U_{m_{i}}\cap U_{n_{i}}=\emptyset$. Then $Y=\{x_{i}:i\in\mathbb{N}\}$ is an effectively discrete subspace of $X$ since $U_{n_{i}}\cap Y=\{x_{i}\}$ for all $i$.
Let $x_{0}$ be any element of $X$ different from $p$, and let $n_{0}$ and $m_{0}$ be such that $e(p,x_{0})=\langle m_{0},n_{0}\rangle$. Now fix $s$, and suppose we have defined $x_{i}$, $n_{i}$, and $m_{i}$ for all $i\leq s$. Since $p\in\bigcap_{i\leq s}U_{m_{i}}$ we have that $p\in U_{\overline{k}(p,\{m_{i}:i\leq s\})}\subseteq\bigcap_{i\leq s}U_{m_{i}}$. And since $p$ is a limit point, we can let $x_{s+1}$ be the least element of $U_{\overline{k}(p,\{m_{i}:i\leq s\})}$ larger than $x_{s}$ and different from $p$. Say $e(p,x_{s+1})=\langle m,n\rangle$. Then $p\in\bigcap_{i\leq s}U_{m_{i}}\cap U_{m}$ and $x_{s+1}\in U_{m_{s}}\cap U_{n}$, so $p\in U_{\overline{k}(p,\{m_{i}:i\leq s\}\cup\{m\})}\subseteq\bigcap_{i\leq s}U_{m_{i}}\cap U_{m}$, and $x_{s+1}\in U_{\overline{k}(x_{i},\{m_{s},n\})}\subseteq U_{m_{s}}\cap U_{n}$. Let $m_{s+1}=\overline{k}(p,\{m_{i}:i\leq s\}\cup\{m\})$ and $n_{s+1}=\overline{k}(x_{s+1},\{m_{s},n\})$.
So, $x_{s+1}\in U_{n_{s+1}}$, $p\in U_{m_{s+1}}$, $U_{m_{s+1}}\subseteq U_{m_{s}}$, and $U_{m_{s+1}}\cap U_{n_{s+1}}=\emptyset$. Fix any $i\leq s$. We have $x_{s+1}\in U_{n_{s+1}}\subseteq U_{m_{s}}$, while $x_{i}\in U_{n_{i}}$ which is disjoint from $U_{m_{i}}$ and hence from $U_{m_{s}}\subseteq U_{m_{i}}$. Thus, $x_{i}\notin U_{n_{s+1}}$ and $x_{s+1}\notin U_{n_{i}}$, as needed. ∎
@@
We construct a computable sequence $\langle V_{n}:n\in\omega\rangle$ of subsets of $\omega$ and then generate a computable CSC space $\langle\omega,\mathcal{U},k\rangle$ using Proposition 3.3. Initially, we enumerate $x$ into $V_{2\langle x,y\rangle}$ and $V_{2\langle x,y\rangle+1}$ for all $y$. We then enumerate more elements into the $V_{n}$ in stages, maintaining throughout that $V_{2\langle x,y\rangle}\cap V_{2\langle y,x\rangle}=\emptyset$ for all $x\neq y$. By Proposition 3.3 (2), we can compute the indices $m$ and $n$ of $V_{2\langle x,y\rangle}$ and $V_{2\langle y,x\rangle}$ in $\mathcal{U}$ uniformly computably from $x$ and $y$. Hence, the function $e:X\times X\to\mathbb{N}$ mapping $\langle x,y\rangle$ to $\langle m,n\rangle$ is computable. Since $x\in V_{2\langle x,y\rangle}$ and $y\in V_{2\langle y,x\rangle}$, this $e$ will witness that $\langle\omega,\mathcal{U},k\rangle$ is effectively Hausdorff.
We aim to additionally satisfy the following requirements for all $x,e,u\in\omega$:
@@
$\mathcal{R}_{e,u}$ ::: if $\Phi_{e}$ is the characteristic function of an infinite set $Y$ and $\Phi_{u}$ is total on $Y$, then there exists $x\in Y$ and $n_{0},\ldots,n_{r}\in\omega$ such that $\Phi_{u}(x)\operatorname{\downarrow}=\langle n_{0},\ldots,n_{r}\rangle$ but $\bigcap_{i\leq r}V_{n_{i}}\neq\{x\}$.
We call these the $\mathcal{D}$-requirements and $\mathcal{R}$-requirements for short. The $\mathcal{D}$-requirements thus ensure that our CSC space is discrete, while the $\mathcal{R}$-requirements ensure that it has no infinite computable effectively discrete subspace. At each stage, at most one requirement will act. If a $\mathcal{D}$-requirement acts it claims an index $n\in\omega$, and if an $\mathcal{R}$-requirement acts it claims finitely many indices $n_{0},\ldots,n_{r}\in\omega$. A requirement may have its claims cancelled at a later stage, and then make new claims later still, as we describe below.
We proceed to the construction. We assign a priority ordering to the requirements in order type $\omega$, as usual. For notational convenience, for $n=2\langle x,y\rangle$ we let $\overline{n}$ denote $2\langle y,x\rangle$. Suppose we are at stage $s\in\omega$. Say $\mathcal{D}_{x}$ requires attention at stage $s$ if it does not currently have a claim. Say $\mathcal{R}_{e,u}$ requires attention at stage $s$ if it does not currently have any claims, and if there exist $x<s$ and $\langle n_{0},\ldots,n_{r}\rangle<s$ as follows:
@@
  • no $n_{i}$ is claimed by any stronger $\mathcal{D}$ requirement;
@@
If it exists, fix the strongest requirement with index smaller than $s$ that requires attention at stage $s$.
If this is $\mathcal{D}_{x}$ for some $x$, choose the least $y$ so that $12\langle x,y\rangle+$ 1 is larger than all indices claimed by any requirement at any point in the construction, including claims that may since have been cancelled. Since at most one requirement claims at most finitely many numbers at any stage, $y$ exists. Now say that $\mathcal{D}_{x}$ claims $12\langle x,y\rangle+$ 1. If the requirement is instead $\mathcal{R}_{e,u}$ for some $e$ and $u$, fix the least witnessing $x,\langle n_{0},\ldots,n_{r}\rangle<s$. Cancel the claims of any lower priority requirements, and say that $\mathcal{R}_{e,u}$ claims $n_{0},\ldots,n_{r}$. In either case, say that the requirement in question acts. Finally, for each $n$ that is claimed by some $\mathcal{R}$-requirement at this point, enumerate $s$ into $V_{n}$. This completes the construction.
Since the only time a number $s$ is enumerated into some $V_{n}$ during the construction is at stage $s$, the sequence $\langle V_{n}:n\in\omega\rangle$ is computable. It follows that the CSC space $\langle\omega,\mathcal{U},k\rangle$ generated from $\langle V_{n}:n\in\omega\rangle$ is computable as well.
Next, note that if $n$ is even and $n\neq\overline{n}$, then $n$ and $\overline{n}$ can never be claimed at the same time (whether by the same requirement, or by different ones). Indeed, even indices $n$ can only be claimed by $\mathcal{R}$-requirements. When an even $n$ is claimed by a new $\mathcal{R}$-requirement at some stage $s$, then $\overline{n}$ must not already be claimed by any higher priority requirement (necessarily another $\mathcal{R}$-requirement). And if $\overline{n}$ was already claimed by a lower priority requirement, then this claim is first cancelled. It follows that, at the end of each stage $s$, the number $s$ is enumerated into at most one of $V_{n}$ or $V_{\overline{n}}$ for every even $n$ with $n\neq\overline{n}$. In particular, if $x\neq y$ then $V_{2\langle x,y\rangle}$ and $V_{2\langle y,x\rangle}$ are disjoint. By our remarks above, $\langle\omega,\mathcal{U},k\rangle$ is effectively Hausdorff.
An induction on the priority ordering shows that each requirement acts only finitely often. This is because once a requirement acts at a stage after all stronger $\mathcal{R}$-requirements have stopped acting, it claims some indices, and these claims cannot later be cancelled. Hence, the requirement can never require attention and act again.
It thus remains only to show that all requirements are satisfied. Fix a requirement, and let $s_{0}$ be least such that every stronger requirement has stopped acting prior to stage $s_{0}$. Thus, either $s_{0}=0$ or some stronger $\mathcal{R}$-requirement acted at stage $s_{0}-1$ and thereby cancelled any claim the requirement in question may have had. Either way, then, this requirement has no claims at the start of stage $s_{0}$. Hence, if the requirement in question is $\mathcal{D}_{x}$ for some $x$, then it requires attention at the least stage $s\geq s_{0}$ with $s>x$, at which point it claims some $12\langle x,y\rangle+$ 1 that has never been claimed by any requirement before. By choice of $s_{0}$, this claim is never cancelled at any stage after $s$, and so in particular, $12\langle x,y\rangle+$ 1 is never claimed by any $\mathcal{R}$-requirement. It follows that no numbers are enumerated into $V_{2\langle x,y\rangle+1}$ during the construction, so $V_{2\langle x,y\rangle+1}$ is just the singleton $\{x\}$ and $\mathcal{D}_{x}$ is satisfied.